### 15 Dec 2023 v1

# GSVA: Generalized Segmentation via Multimodal Large Language Models

Abstract
Generalized Referring Expression Segmentation (GRES) extends the scope of classic RES to referring to multiple objects in one expression or identifying the empty targets absent in the image. GRES poses challenges in modeling the complex spatial relationships of the instances in the image and identifying non-existing referents. Recently, Multimodal Large Language Models (MLLMs) have shown tremendous progress in these complicated vision-language tasks. Connecting Large Language Models (LLMs) and vision models, MLLMs are proficient in understanding contexts with visual inputs. Among them, LISA, as a representative, adopts a special [SEG] token to prompt a segmentation mask decoder, e.g., SAM, to enable MLLMs in the RES task. However, existing solutions to of GRES remain unsatisfactory since current segmentation MLLMs cannot properly handle the cases where users might reference multiple subjects in a singular prompt or provide descriptions incongruent with any image target. In this paper, we propose Generalized Segmentation Vision Assistant (GSVA) to address this gap. Specifically, GSVA reuses the [SEG] token to prompt the segmentation model towards supporting multiple mask references simultaneously and innovatively learns to generate a [REJ] token to reject the null targets explicitly. Experiments validate GSVA’s efficacy in resolving the GRES issue, marking a notable enhancement and setting a new record on the GRES benchmark gRefCOCO dataset. GSVA also proves effective across various classic referring expression segmentation and comprehension tasks. Code will be available at https://github.com/LeapLabTHU/GSVA.
1 Introduction
Figure 1: Comparison of the segmentation masks by LISA [24] and GSVA, facing the challenges in Generalized Referring Expression Segmentation (GRES) [30]. (a) LISA fails to segment the correct targets when multiple targets are requested due to the single [SEG] token restriction. GSVA successfully generates all target masks via learning multiple [SEG] tokens. (b) When the referent does not exist in the image, i.e., the empty target is requested, LISA reluctantly produces the wrong mask because of the compulsive [SEG] token output. In contrast, GSVA can reject the empty targets by predicting [REJ] tokens in the output sequence.
Referring Expression Segmentation (RES) [7, 18] is an emerging vision-language (VL) task that aims to predict the masks of the corresponding objects referred to in the language expression. RES has great potential in many areas, especially embodied AI [57, 43, 16, 49], including VL navigation, human-robot interaction, etc. Nevertheless, the simplification in RES formulation that one referring expression must match an individual object in the image [22] has left a gap between current RES algorithms and real-world applications, neglecting multiple-target and empty-target cases.
To bridge this gap, the Generalized Referring Expression Segmentation (GRES) [30] task has recently been proposed to support multiple-target and empty-target cases. In practical scenarios, users tend to refer to multiple subjects within a single instruction or provide descriptions that do not correspond to any targets in the image. As an example shown in Figure 1 (a), the left man and the second man from the right are targeted simultaneously, while a banana is referred to in a scene of apples in Figure 1 (b). RES takes no account of these cases, which are handled by GRES. In addition to the multimodal correspondences between images and text prompts in classic RES, GRES poses new challenges in handling more complicated multiple-target and empty-target cases. Therefore, the models are required to handle complex spatial relationships of the instances in the image [30] to segment the targets at various locations and reject the empty targets in the wrong places.
The recent blooming Multimodal Large Language Models (MLLMs) [1, 68, 64, 31, 26] meet the requirements of GRES since they show excellency in complex reasoning [9] and instruction following [41] with visual inputs by aligning the LLMs [2, 40, 52, 53, 51, 8] and Visual Foundation Models (VFMs) [44, 54, 14, 60]. To support segmentation, LISA [24], a representative MLLM, incorporates LLaVA [31] and SAM [23] by a special [SEG] token as the prompt query, bridging the MLLM and the Segmentation Foundation Model (SFM) to generate masks of the referents in the user’s instructions. Successfully though LISA handles RES, GRES is still beyond its reach. As shown in Figure 1, LISA fails to work well in GRES where the multiple-target and empty-target challenges remain uncharted.
To address the above challenges, we propose Generalized Segmentation Vision Assistant (GSVA). We attribute the vulnerability of LISA in GRES to the single permanent [SEG] token that restricts its flexibility. Therefore, we present two pivotal designs in GSVA: (1) learning to predict multiple [SEG] tokens to segment multiple targets; (2) rejecting empty targets in referring expressions by predicting [REJ] tokens. Specifically, when multiple targets are requested in the referring expression, we change the output sequence from a single [SEG] token to multiple [SEG] tokens corresponding to the entities in the expression, which encourages the MLLM to learn to output multiple [SEG] tokens. To distinguish each [SEG] token and avoid ambiguity, we add the expression of each entity in front of the corresponding [SEG] token, which can also be regarded as implicit In-Context Learning, hinting each [SEG] token to focus on the specific target. Meanwhile, if referents are absent in the image, the corresponding [SEG] tokens after the prompts are altered to [REJ] tokens to identify empty targets. The predicted [REJ] tokens are directly assigned with empty masks without being fed into the SFM, which liberates the segmentation model from seeking non-existing targets in the image. Benefiting from these novel designs, GSVA takes a big step forward in addressing GRES challenges, as shown in the second row of Figure 1.
In summary, our contributions are as follows: we propose two novel designs for segmentation MLLM to address the GRES challenges, including learning multiple [SEG] tokens along with the referring expressions and rejecting empty targets via [REJ] tokens. Extensive experiments demonstrate the effectiveness of the proposed GSVA. On the GRES benchmark gRefCOCO [30], GSVA outperforms LISA [24] by large margins, also surpassing previous SOTA methods, which sets new records with over 70% gIoU and cIoU. In addition, extensive experiments in RefCOCO, RefCOCO+ [22], and RefCOCOg [37] demonstrate that GSVA is also competitive on the classic RES and Referring Expression Comprehension (REC) tasks, achieving nearly 80% cIoU for RES and 90% Prec@0.5 for REC.
2 Related Works
RES and GRES. Referring Expression Segmentation (RES) [7, 18, 6, 37] aims to segment targets in the image based on language expressions. Under the assumption that one expression matches one existing target, many works explore fusing image and language [29, 27, 4, 15, 62]. Currently, most RES methods adopt the cross-attention module to bridge the modality gap [48, 63, 70, 59]. Another line of research enables the text prompts for segmentation model with a unified decoder architecture, including SAM [23], X-Decoder [75], SEEM [76], and Grounded-SAM which combines GroundingDINO [33] with SAM, etc. Although these works open a brilliant chapter for RES, the one-to-one restriction still holds. To break the jail for the arbitrary number of targets, DMMI [19] focuses on the one-to-many setting where text expression refers to varying numbers of targets. ReLA [30] proposes the Generalized Referring Expression Segmentation (GRES) task, supporting both the multi-target and empty-target scenarios, which is the main scope of our research of GSVA.
MLLM. Multimodal Large Language Models (MLLMs) align the vision and language modalities by various techniques, including cross-attention module [1], prompt tuning tokens [68], Q-Former [26, 10], linear projection layers [31], and unified model architectures [42, 60]. Endowed with unprecedented capabilities in reasoning with world-knowledge and following instructions of users, MLLM shows extraordinary performances in various vision language tasks [36, 56]. Equipped with the segmentation decoders or detection heads, MLLMs can also excel in the vision-centric tasks, such as object detection and segmentation [55, 67, 61, 47, 28]. Among them, LISA [24] makes the most of the reasoning ability with a [SEG] token to address the Reasoning Segmentation problem. However, LISA fails to tackle the challenge in GRES due to the inflexible [SEG] token, which is addressed by our proposed GSVA.
3 Generalized Segmentation Vision Assistant
Figure 2: Overview of GSVA. At the bottom of the figure, the MLLM encodes the input image and concatenates the tokenized text tokens to follow instructions. GSVA generates multiple [SEG] tokens to handle multiple referred targets and rejects the objects absent in the image through [REJ] tokens. At the top of the figure, the SFM also encodes the image for segmentation and selects all [SEG] tokens in the output sequence to prompt the mask decoder to segment the target objects referred to in the instructions.
In this section, we initiate with the introduction of the model design of Generalized Segmentation Vision Assistant (GSVA), which is followed by the analysis of some certain limitations of LISA in Generalized Referring Segmentation (GRES). Subsequently, we delve into the introduction of two pivotal elements of GSVA, segmenting multiple targets and learning the rejection token. These components are fundamental in the conceptualization and construction of GSVA.
3.1 Model Architecture
The architecture of GSVA is illustrated in Figure 2, resembling LISA [24], which enables high-fidelity segmentation outputs by integrating two types of foundation models: (1) Multimodal Large Language Model (MLLM) as an aligned vision-language cognitive module; (2) Segmentation Foundation Model (SFM) to segment the target out of the input image based on user’s instruction. To connect these two modules, LISA proposes a paradigm named embedding as mask where an extra [SEG] token is appended to the vocabularies of the MLLM and serves as the prompt of the SFM to segment the target following the intention of the user.
Multimodal Large Language Model. The MLLM consists of a decoder-based language model $F_{\text{LLM}}$ to auto-regressively generate text responses following the user’s inputs, a vision encoder $F_{\text{V1}}$ to extract features from the input image, and a linear projector $\phi$ to align the representations between image and text modalities. Specifically, the pretrained LLaVA [31] variants with CLIP-ViT-L/14 [44] and Vicuna-7B/13B [8] are employed. Given an input image $\mathbf{x}_{\text{img}}$, the vision encoder $F_{\text{V1}}$ first encodes it into image features, and then the projector $\phi$ maps the features into the visual token embeddings in the LLM input space:
$\mathbf{h}_{\text{img}}=\phi(F_{\text{V1}}(\mathbf{x}_{\text{img}})),$ (1)
where the input image $\mathbf{x}_{\text{img}}$ is typically resized to $h\!\times{}\!w\!\times{}\!3$, and the image tokens $\mathbf{h}_{\text{img}}\in\mathbb{R}^{n_{\text{img}}\!\times{}\!d}$ is aligned with the language modality. For CLIP-ViT-L/14, the input image with $h\!=\!w\!=\!224$ is encoded with ViT of patch size in 14, therefore the length of tokens $n_{\text{img}}\!=\!hw/14^{2}\!=\!256$, and the LLM dimensions $d$ are 4096 and 5120 for Vicuna-7B/13B, respectively. Along with the input image, the text instructions describing the targets to segment are tokenized into text tokens by the LLM tokenizer $T$:
$\mathbf{h}_{\text{txt}}=T(\mathbf{x}_{\text{txt}}).$ (2)
The image tokens and text tokens are concatenated together and then fed into the LLM after prepending a sequence of fixed prompt tokens $\mathbf{h}_{\text{prompt}}$ (omitted in the figure). The output token embeddings $\tilde{\mathbf{y}}_{\text{txt}}$ are generated auto-regressively:
$\tilde{\mathbf{y}}_{\text{txt}}=F_{\text{LLM}}\left([\mathbf{h}_{\text{prompt}}||\mathbf{h}_{\text{img}}||\mathbf{h}_{\text{txt}}]\right),$ (3)
where ||||| | is concatenation operation. The text responses are obtained from $\tilde{\mathbf{y}}_{\text{txt}}$ by applying a linear classifier to predict the next words in the vocabulary.
In LISA, a special token [SEG] is appended in the vocabulary to activate the segmentation ability of MLLM. The model learns to predict [SEG] token in the output sequence to indicate there is a target to segment. LISA then selects the output embedding of the [SEG] token $\tilde{\mathbf{y}}_{\text{txt}}[\text{SEG}]$, and projects it into the prompt space of the SFM by an MLP projector $\psi$:
$\mathbf{h}_{\text{seg}}=\psi\left(\tilde{\mathbf{y}}_{\text{txt}}[\text{SEG}]\right).$ (4)
The segmentation model is hence ready to decode the target mask from the query token $\mathbf{h}_{\text{seg}}$.
Segmentation Foundation Model. The SFM is a query-based segmentation model, where a frozen vision encoder $F_{\text{V2}}$ takes in images of higher resolution than the vision encoder in MLLM to keep more details, followed by a trainable mask decoder $F_{\text{mask}}$ to decode masks from the queries. The pretrained SAM [23] with ViT-H backbone is instantiated as the SFM to produce high-quality masks. The given input image $\mathbf{x}_{\text{img}}$ is resized to a larger resolution $H\!\times{}\!W\!\times{}\!3$, with $H\!=\!W\!=\!1024$ and encoded with $F_{\text{V2}}$ to extract features $\mathbf{f}_{\text{seg}}$ for segmentation:
$\mathbf{f}_{\text{seg}}=F_{\text{V2}}(\mathbf{x}_{\text{img}}).$ (5)
Condition on the features $\mathbf{f}_{\text{seg}}\in\mathbb{R}^{\frac{H}{16}\!\times{}\!\frac{W}{16}\!\times{}\!C}$ with $C\!=\!256$ for SAM, the mask decoder $F_{\text{mask}}$ decodes the segmentation masks from the segmentation queries $\mathbf{h}_{\text{seg}}\in\mathbb{R}^{N_{\text{seg}}\times{}C}$:
$\tilde{\mathbf{y}}_{\text{mask}}=F_{\text{mask}}(\mathbf{h}_{\text{seg}}|\mathbf{f}_{\text{seg}}),$ (6)
where each query in $\mathbf{h}_{\text{seg}}$ corresponds to one segmentation mask in $\tilde{\mathbf{y}}_{\text{mask}}$.
LISA assumes that only one target exists to segment in the input image and its corresponding instructions. However, in GSVA, we extend it to a new scenario with multiple targets and empty targets, including multiple [SEG] tokens to invoke segmentation and [REJ] tokens to reject unreasonable instructed targets absent in the image. As shown in Figure 2, GSVA supports multiple [SEG]/[REJ] tokens in the output sequence and selects all the [SEG] tokens and discards every [REJ] token after attaining $\tilde{\mathbf{y}}_{\text{txt}}$ in Eq. (3). Therefore, there is more than one query in $\mathbf{h}_{\text{seg}}$, thus enabling the SFM to segment multiple targets. These designs make GSVA competent in the GRES task, where the awareness of multiple and empty targets is of vital importance.
3.2 GRES: Task and Challenges
[…]
Challenges. The challenges in GRES are common in practice, especially in embodied AI [57, 43, 16, 49]. The one challenge is multiple targets. Take robot navigation and planning [39, 57, 6] as an example, a robot may be asked to perceive multiple targets in the surrounding environment, e.g., to bring the two bowls of soup in Figure 2. The vision system of the robot needs to locate and segment the containers holding the referred food one at a time. The other challenge is empty target. Suppose the robot is ordered to cut an apple with a knife in the scene of Figure 2, whereas no apple is in the view of the camera, the vision system has to identify that the referred object is not in the scene. If it relies on some conventional RES methods which assume the expression must match something in the image, the output of the vision system could be undefined and potentially dangerous in some real-world cases.
Differences from ReasonSeg. The Reasoning Segmentation (ReasonSeg) task is proposed by LISA [24], which emphasizes the complex text instructions in RES. In ReasonSeg, the instructions are more implicit and sophisticated than RES, including common sense reasoning and world knowledge. The logic chain is usually longer and more challenging in ReasonSeg, which requires the model to reason from the implicit query texts and deduce the target object referred to in the image. Nevertheless, GRES introduces complexity to RES in another dimension by involving complicated spatial relationships of objects. Hence, the model has to learn this spatial complexity and be capable of understanding the relationships between these instances. To meet these requirements, LISA tunes MLLMs with logically complex instructions paired with masks. In contrast, our proposed GSVA arouses the spatial modeling capabilities of MLLMs by learning multiple targets and rejecting empty targets. These designs are elaborated as follows.
3.3 Multiple [SEG] Tokens for Multiple Targets
Single [SEG] token. LISA [24] follows the classic RES methods to generate a segmentation mask under given instructions by adding the [SEG] token into the answer. The prompt is formatted as:
User: What is {obj} in this image? Please output segmentation mask.
Assistant: Sure, it is [SEG].
In the above prompt, {obj} represents an instance referred to or some semantic area to segment, and the embeddings of the [SEG] token output from the LLM are projected to prompt the SFM. When multiple instances are requested simultaneously, this prompt would confuse the model since only one [SEG] token is forced to match several targets. As shown in Figure 1 (a), LISA coercively predicts the masks of the left man and the guy second from the right, leading to a tattered mask and masking the wrong location.
Figure 3: Example of the prompts and predicted masks of GSVA-Vicuna-7B drawn from gRefCOCO validation set. (a) depicts the multiple-target case, in which two zebras referred to are handled with two separate [SEG] tokens. (b) shows the empty-target case, where no apple is in the bowl. Thus, the null referent is rejected with a [REJ] token, and no segmentation mask will be generated.
Multiple [SEG] tokens. To mitigate this issue in GRES, we relax this constraint for GSVA to support multiple target outputs via learning multiple [SEG] tokens. To avoid the ambiguity between the [SEG] tokens and the corresponding objects, we prepend the referring expression before each [SEG] token, i.e., each segmentation prompt in the output text is {obj_n}:[SEG], as shown in Figure 3 (a). Streaming all the pairs of object and [SEG] token into one sentence, the question-answer prompt is formatted as (e.g., two targets):
User: What are {obj1}, {obj2} in this image? Please output the segmentation masks. Assistant: Sure, {obj1}:[SEG], {obj2}:[SEG].
This prompt requests the MLLM to identify and distinguish different objects in the image based on the instructions and infuse the corresponding location information of each target into the associated [SEG] token. We regard this ability as an implicit multimodal version of In-Context Learning (ICL), which is demonstrated by many prior works [2, 12, 71, 50, 25]. GSVA takes the target description preceding each [SEG] token as a hint to link this token to the object requested in the image through auto-regressive decoding.
3.4 Rejecting Empty Targets via [REJ] Tokens
Empty targets. In GRES, many expressions match no targets in the image, including absence, incorrect attributes, inaccurate locations, etc. These expressions should be treated as empty targets for models to predict all-negative masks. LISA [24] falls short of predicting masks with all-zeros seamlessly since the [SEG] token always calls for a segmentation mask. As shown in Figure 1 (b), LISA incorrectly marks a piece of green apple as the empty target banana.
[REJ] token. We let the MLLM of GSVA predict a [REJ] token for each object that does not exist in the image but is referred to in the instructions, as shown in Figure 3 (b). GSVA predicts the targets marked with [REJ] tokens as empty targets, therefore setting all-zero masks for them. The involvement of [REJ] tokens directly rejects the empty target, liberating the mask decoder of the SFM from learning to segment the inexistent targets. An example prompt with one existing target and one empty target is as follows:
User: What are {obj1} (does not exist), {obj2} in this image? Please output the segmentation masks. Assistant: Sure, {obj1}:[REJ], {obj2}:[SEG].
The [REJ] token prediction can also be seen as a variant of VQA task, where the specified object and its position in the image need to be considered. Thanks to the unprecedented capabilities of MLLM in understanding the images [31, 74, 26, 10] and reasoning the spatial relationships of the referring objects [42, 69, 72], we make the most of the MLLM in GSVA to unleash the burden of the segmentation model. The proposed empty-target-aware mechanism both improves the quality of masks and ameliorates the errors of identifying nonexistent objects.
4 Experiments
In this section, we conduct comprehensive experiments to validate the efficacy of GSVA. First, we show the results on gRefCOCO [30] dataset to show the superiority of GSVA in GRES tasks. Then we verify GSVA that is also competent with other baselines in classic RES, REC tasks. We move on to ablate some important design choices of GSVA, followed by some qualitative visualization of GRES results.
Generalized Referring Expression Segmentation on gRefCOCO
Method Validation Set Test Set A Test Set B
gIoU cIoU N-acc. gIoU cIoU N-acc. gIoU cIoU N-acc.
MattNet [66] 48.24 47.51 41.15 59.30 58.66 44.04 46.14 45.33 41.32
LTS [21] 52.70 52.30 - 62.64 61.87 - 50.42 49.96 -
VLT [11] 52.00 52.51 47.17 63.20 62.19 48.74 50.88 50.52 47.82
CRIS [58] 56.27 55.34 - 63.42 63.82 - 51.79 51.04 -
LAVT [62] 58.40 57.64 49.32 65.90 65.32 49.25 55.83 55.04 48.46
ReLA [30] 63.60 62.42 56.37 70.03 69.26 59.02 61.02 59.88 58.40
LISA-Vicuna-7B [24] 32.21 38.72 2.71 48.54 52.55 6.37 39.65 44.79 5.00
GSVA-Vicuna-7B 63.32 61.70 56.45 70.11 69.23 63.50 61.34 60.26 58.42
LISA-Vicuna-7B [24] (ft) 61.63 61.76 54.67 66.27 68.50 50.01 58.84 60.63 51.91
GSVA-Vicuna-7B (ft) 66.47 63.29 62.43 71.08 69.93 65.31 62.23 60.47 60.56
LISA-Vicuna-13B [24] 32.73 39.85 3.66 48.76 53.62 4.89 39.49 45.35 4.41
GSVA-Vicuna-13B 62.97 60.18 58.44 67.17 67.59 54.60 58.06 57.28 52.22
LISA-Vicuna-13B [24] (ft) 63.45 62.99 55.25 68.18 69.65 52.16 61.84 62.24 56.15
GSVA-Vicuna-13B (ft) 68.01 64.05 65.36 71.75 70.51 67.25 63.83 61.28 63.11
LISA-Llama2-13B [24] 33.26 39.64 3.27 49.76 53.80 7.28 40.49 45.41 5.73
GSVA-Llama2-13B 63.20 62.38 54.51 69.52 69.86 57.84 62.06 60.77 58.30
LISA-Llama2-13B [24] (ft) 65.24 63.96 57.49 69.99 71.00 55.43 62.11 62.29 56.34
GSVA-Llama2-13B (ft) 70.04 66.38 66.02 73.29 72.79 64.72 65.45 63.20 62.47
Table 1: Generalized referring expression segmentation (GRES) results on gRefCOCO [30] dataset. gIoU and cIoU are IoU metrics averaged by each example and accumulated over whole dataset, respectively. N-acc. is short for the accuracy of correctly classifying null targets. Baselines are copied from Liu et al. [30]. (ft) denotes the model is finetuned on the training set of gRefCOCO.
4.1 GRES
Settings. We adopt gRefCOCO [30] dataset to validate GSVA and LISA [24] on GRES, which contains 278,232 expressions, including 80,022 multi-target and 32,202 empty-target ones, referring to the objects in 19,994 images. The images are split into four subsets: training, validation, test-A, and test-B, following the same UNC partition of RefCOCO [65]. We first add the gRefCOCO training set into the mixed training dataset in LISA to pretrain GSVA and LISA for 50,000 steps and then finetune the models on the gRefCOCO training dataset for another 10 epochs. We evaluate the pretrained and finetuned models on the remaining validation set, test set A, and test set B, respectively. We adopt the gIoU, cIoU metrics for the segmentation mask outputs. Following the implementation in Liu et al. [30], gIoU averages the IoU for each mask, whereas cIoU computes the cumulative intersection area over the cumulative union area across the whole dataset. As for the empty target, we compute the No-target-accuracy (N-acc.), which is the ratio of the correctly classified empty-target expressions over all the empty-target expressions in the dataset. For a correctly classified empty target, the gIoU is set to 1.0, and the cIoU does not take them into account, while for a misclassified empty target, the gIoU is set to 0.0 and the union area is accumulated in the cIoU. Following the criteria in Liu et al. [30], a predicted mask is regarded as empty for LISA if the positive pixels are less than 50, whereas GSVA predicts [REJ] tokens to identify empty targets.
Results. We report the GRES segmentation results of GSVA and LISA [24] in Table 1. Three variants of GSVA, including GSVA-Vicuna-7B, GSVA-Vicuna-13B, and GSVA-Llama2-13B show competitive performance without finetuning on gIoU to the strongest non-LLM baseline ReLA [30]. However, LISA models fail to handle the GRES task without finetuning, showing degradation in both gIoU and cIoU in each model variant. Especially the low N-acc indicates that LISA is unable to correctly reject the empty targets. When finetuned on gRefCOCO training set, GSVA-Vicuna-7B performs better than the finetuned LISA counterpart, with about 4% improvement in gIoU and over 5% in N-acc on all three evaluation splits. GSVA variants with larger LLM incorporated further push the limits, achieving over 70% in gIoU on the validation set, 73% on test set A, and 65% on test set B. The 13B models also consistently outperform LISA by large margins, demonstrating the superiority of GSVA in GRES task.
4.2 Referring Expression Segmentation
Settings. To validate the ablitiies to handle various tasks, we evaluate GSVA in the classic RES task. Following the common evaluation protocols, we test variants of GSVA and LISA equipped with different LLMs on RefCOCO, RefCOCO+ [22], and RefCOCOg [37]. We follow the UNC split to perform experiments on RefCOCO and RefCOCO+, and UMD split for RefCOCOg. The models are firstly pretrained as in GRES, and then finetuned for 10 epochs with a joint dataset of these three RES training set. The cIoU metric is adopted to measure the model performances.
Results. Table 2 shows the RES results of GSVA. For the pretrained models, all the three variants of GSVA achieve higher cIoU than LISA [24] by clear margins. Our 7B model outperforms LISA-Vicuna about 2% cIoU on almost every data split. After finetuning, the preponderance of GSVA over LISA keeps and even enlarges. Specifically, the cIoU metrics of GSVA-Llama2-13B on 8 sets surpass LISA by at least 2.8%. For the test set of RefCOCOg, the margin even grows to 5.9%, which exhibits GSVA is also competitive in the classic RES task.
Referring Expression Segmentation on RefCOCO, RefCOCO+, and RefCOCOg
Method RefCOCO (UNC) RefCOCO+ (UNC) RefCOCOg (UMD)
Val. Test-A Test-B Val. Test-A Test-B Val. Test
MCN [35] 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
VLT [11] 67.5 70.5 65.2 56.3 61.0 50.1 55.0 57.7
CRIS [58] 70.5 73.2 66.1 62.3 68.1 53.7 59.9 60.4
LAVT [62] 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
ReLA [30] 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
X-Decoder [75] - - - - - - 64.6 -
SEEM [76] - - - - - - 65.7 -
PolyFormer-L [32] 76.0 78.3 73.3 69.3 74.6 61.9 69.2 70.2
$*{}^{*}$ [24] 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5
GSVA-Vicuna-7B 76.4 77.4 72.8 64.5 67.7 58.6 71.1 72.0
$*{}^{*}$ [24] (ft) 74.9 79.1 72.3 65.1 70.8 58.1 67.9 70.6
GSVA-Vicuna-7B (ft) 77.2 78.9 73.5 65.9 69.6 59.8 72.7 73.3
LISA-Vicuna-13B [24] 71.7 74.7 68.1 59.4 64.2 52.9 65.2 66.1
GSVA-Vicuna-13B 74.6 77.5 70.5 62.5 66.5 55.5 69.6 71.2
LISA-Vicuna-13B [24] (ft) 76.0 78.8 72.9 65.0 70.2 58.1 69.5 70.5
GSVA-Vicuna-13B (ft) 78.2 80.4 74.2 67.4 71.5 60.9 74.2 75.6
LISA-Llama2-13B [24] 73.4 76.2 69.5 62.3 66.6 56.3 68.2 68.5
GSVA-Llama2-13B 77.7 79.9 74.2 68.0 71.5 61.5 73.2 73.9
LISA-Llama2-13B [24] (ft) 76.3 78.7 72.4 66.2 71.0 59.3 70.1 71.1
GSVA-Llama2-13B (ft) 79.2 81.7 77.1 70.3 73.8 63.6 75.7 77.0
Table 2: Referring expression segmentation results on RefCOCO, RefCOCO+ [22] and RefCOCOg [37] dataset. The cIoU metrics of each split are reported. Baselines are excerpted from Lai et al. [24]. (ft) denotes the models are finetuned on the joint training set of the referring expression segmentation datasets. * means the results are excerpted from the original paper.
4.3 Referring Expression Comprehension
Settings. Since GSVA is capable for RES tasks, it is natural to transfer to Referring Expression Comprehension (REC) tasks, by simply computing the bounding boxes of the masks. The datasets of REC are the same as RES, in which we evaluate GSVA. If the IoU of a predicted bounding box and the ground truth is greater than 0.5, this prediction is marked as correct. We use the same models in RES to evaluate the phrase grounding capability in REC tasks.
Results. As shown in Table 3, we mainly compare our GSVA to LISA in the three variants. Without finetuning, GSVA-Vicuna-7B outperforms LISA with a large margin over 7% in almost all evaluation sets. The similar trends also hold for the Vicuna-13B and Llama2-13B variants. GSVA also benefits a lot from finetuning, e.g., the 7B variant achieves consistenly higher Prec@0.5 than uLLaVA-7B [61] with LoRA finetuning, which is another strong baseline that adopts the “mask2bbox” pipelines without direct bounding box supervision. The finetuned GSVA also shows competitive performances to the fully finetuned uLLaVA-7B, suggesting the strong potential of our method. With larger LLMs incorporated, the performances of GSVA continue with over 3% increments over all datasets.
Referring Expression Comprehension on RefCOCO, RefCOCO+, and RefCOCOg
Method RefCOCO RefCOCO+ RefCOCOg
Val. Test-A Test-B Val. Test-A Test-B Val. Test
u-LLaVA-7B [61] (LoRA) 83.47 87.13 80.21 68.74 76.32 60.98 76.19 78.24
u-LLaVA-7B [61] (full-ft) 86.04 89.47 82.26 74.09 81.16 66.61 79.87 81.68
LISA-Vicuna-7B [24] 78.68 81.72 75.74 62.92 68.93 56.49 70.10 72.47
GSVA-Vicuna-7B 85.50 88.01 82.49 70.21 75.62 65.11 79.00 79.21
LISA-Vicuna-7B [24] (ft) 85.39 88.84 82.59 74.23 79.46 68.40 79.34 80.42
GSVA-Vicuna-7B (ft) 86.27 89.22 83.77 72.81 78.78 68.01 81.58 81.83
LISA-Vicuna-13B [24] 80.01 83.26 76.26 63.77 70.24 57.42 71.79 73.34
GSVA-Vicuna-13B 83.12 87.01 80.54 68.14 73.90 62.00 77.08 78.89
LISA-Vicuna-13B [24] (ft) 85.92 89.05 83.16 74.86 81.08 68.87 80.09 81.48
GSVA-Vicuna-13B 87.71 90.49 84.57 76.52 81.69 70.35 83.90 84.85
LISA-Llama2-13B [24] 82.52 85.56 78.82 67.91 73.77 62.25 75.37 76.83
GSVA-Llama2-13B 86.99 89.54 84.08 73.89 79.10 69.38 80.68 82.07
LISA-Llama2-13B [24] (ft) 85.91 88.84 81.73 74.46 80.56 68.26 80.09 81.27
GSVA-Llama2-13B 89.16 92.08 87.17 79.74 84.45 73.41 85.47 86.18
Table 3: Referring expression comprehension results on RefCOCO, RefCOCO+ [22] and RefCOCOg [37] dataset. The metric is precision @ 0.5 IoU threshold. (LoRA) means the LLM in u-LLaVA [61] is finetuned with LoRA adapter, as LISA and GSVA, while (full-ft) represents the LLM in u-LLaVA is fully trained. Results of u-LLaVA-7B with “mask2bbox” strategy are reported for fair comparison.
Figure 4: Visualizations of GSVA and LISA [24] in the GRES task. The first row shows LISA’s segmentation results, the second row is the masks and rejections of GSVA, and the third row shows the referring expressions in the instructions. In (a) multiple target cases, each target is colored with a specific color. In (b) empty target cases, the images turn darker to highlight the incorrect predictions of LISA. The examples are selected from the gRefCOCO validation set. The masks are generated by the 7B models. Zoom in for the best view.
4.4 Ablation Study
Model w/ Vicuna-7B Modifications gRefCOCO Val.
RefExp.
+[SEG]
Multiple
[SEG]
[REJ]
Token
gIoU cIoU N-acc.
GSVA 63.32 61.70 56.45
51.57 60.95 30.32
44.86 59.37 11.96
LISA [24] 32.21 38.72 2.71
21.83 27.22 0.00
Table 4: Ablation study on the core designs of GSVA. ✓ means the employment of the component while ✗ means not. “RefExp.+[SEG]”, “Multiple [SEG]”, and “[REJ] Token” are short for adding referring expression before [SEG] in the answer prompt, using multiple [SEG] tokens, involving [REJ] token, respectively.
[…]
Answers without referring expression. To examine the efficacy of the addition of referring expression in front of [SEG] token, we remove all the referring expressions before the [SEG] tokens. Based on the removal of multiple [SEG]s and [REJ]s, erasing the added referring expression falls back to the original LISA model, as shown in ${}^{\text{th}}$ row of Table 4. To further study the of this addon, we choose only to remove it from GSVA model, keeping other configurations unchanged. Specifically, if there are two referents, the prompts in answer will turn to Assistant: Sure, [SEG], [SEG]. This modification leads to the last row in Table 4. The zero N-acc shows the model cannot learn to identify empty target without the help of the expressions, meanwhile the poor gIoU and cIoU indicates the segmentation ability is damaged since the ambiguity in the anonymous [SEG] tokens. This phenomenon also suggests that the added referring expression hint GSVA to associate each [SEG] token to its corresponding target, which is in coherence with our hypothesis.
4.5 Visualization
We visualize some qualitative results of GSVA to verify its effectiveness. As shown in Figure 4, we present two groups of examples from the validation set of gRefCOCO [30] to see how GSVA outperforms LISA in the face of the two main challenges in GRES: multiple targets and empty targets. In part (a), GSVA has managed to segment all the targets referred to, while LISA could only segment one of the requested instances. For example, in the third column, LISA only predicts the mask of the rightmost vase. On the contrary, GSVA separately segments all three targeted vases. In part (b), LISA mistakenly segments the instances in the image that disagree with the referring expression, e.g., in the sixth column, LISA proposes a mask of a cat in response to the request of the dog on the suitcase, whereas GSVA successfully reject all the empty targets.
5 Conclusion
This paper introduces a novel multimodal large language model dubbed Generalized Segmentation Vision Assistant (GSVA). By introducing multiple [SEG] tokens and the new [REJ] token, GSVA effectively achieves multi-objective segmentation and empty target rejection, which addresses the challenging segmentation problems in practical application scenarios, e.g., Generalized Referring Expression Segmentation (GRES). Extensive experiments on GRES, classic RES, and REC tasks fully demonstrate the superior performance of our method, highlighting its significance for future research and applications.
Acknowledgments
We thank Dr. Yuan Yao and Prof. Zhiyuan Liu for their insightful and valuable comments on this research project. We also appreciate their support of computational resources.
Appendix
Appendix A Implementation Details
We elaborate on the implementation details of GSVA and the settings of the experiments and list all the hyper-parameters in Table 5 for easy reproduction of the results.
Experiment Configuration Model
GSVA-Vicuna-7B
GSVA-Vicuna-13B
GSVA-Llama2-13B
Pretraining Dataset Types SemSeg, RES, VQA, ReasonSeg
SemSeg Datasets ADE20K [73], COCO-Stuff [3], Maplilary Vistas [38], PACO-LVIS [46], Pascal-Part [5]
RES Datasets RefCOCO [22], RefCOCO+ [22], RefCOCOg [37], RefCLEF [22], gRefCOCO [30]
VQA Datasets LLaVA-Instruct-150K [31]
ReasonSeg Datasets ReasonSeg [24]
Epochs / Steps 50,000 steps, gradient accumulation: 10 steps / update
Optimizer AdamW [34], learning rate: $3\!\times{}\!10^{-4}$, weight decay: 0.0, gradient clip: 1.0
ZeRO [45] Stage: 2
[…]
LoRA [17] rank
8
64
64
Finetuning on gRefCOCO [30] Dataset Types RES
SemSeg Datasets -
RES Datasets gRefCOCO [30]
VQA Datasets -
ReasonSeg Datasets -
Epochs / Steps 10 epochs, gradient accumulation: 10 steps / update
Optimizer AdamW [34], learning rate: $1\!\times{}\!10^{-4}$, weight decay: 0.0, gradient clip: 1.0
ZeRO [45] Stage: 2
[…]
LoRA [17] rank
8
64
64
Finetuning on classic RES Dataset Types RES
SemSeg Datasets -
RES Datasets RefCOCO [22], RefCOCO+ [22], RefCOCOg [37], RefCLEF [22]
VQA Datasets -
ReasonSeg Datasets -
Epochs / Steps 50,000 steps, gradient accumulation: 10 steps / update
Optimizer AdamW [34], learning rate: $1\!\times{}\!10^{-4}$, weight decay: 0.0, gradient clip: 1.0
ZeRO [45] Stage: 2
[…]
LoRA [17] rank
8
64
64
Table 5: Detailed configurations and hyper-parameters of GSVA in pretraining and finetuning stages. “steps” mean the model is trained for given steps, which is implemented by fixing the steps in each epoch and sample data from the datasets, while “epochs” mean the model is trained across the whole dataset in each epoch.
Pretraining. In the pretraining stages, GSVA starts from a pretrained MLLM, e.g., LLaVA-Vicuna-7B [31] and a pretrained SFM, e.g., SAM-ViT-H [23]. As shown in Figure 2 in the main paper, the mask decoder $F_{\text{mask}}$, the segmentation query projector $\psi$, the LoRA [17] adapter weights on query and value projections in the LLM, and the token embeddings in the vocabularies are trainable. During training, the cross entropy language modeling loss on the output text sequence $\tilde{\mathbf{y}}_{\text{txt}}$ and the ground truth response $\mathbf{y}_{\text{txt}}$, the output mask $\tilde{\mathbf{y}}_{\text{mask}}$ is supervised by the combination of a binary cross entropy loss and a DICE loss to minimize the error to the ground truth mask $\mathbf{y}_{\text{mask}}$. To summarize, the final objective is
$\mathcal{L}_{\text{total}}$ $=\lambda_{1}\mathcal{L}_{\text{LM}}(\tilde{\mathbf{y}}_{\text{txt}},\mathbf{y}_{\text{txt}})+\lambda_{2}\mathcal{L}_{\text{BCE}}(\tilde{\mathbf{y}}_{\text{mask}},\mathbf{y}_{\text{mask}})$ (7)
$+\lambda_{3}\mathcal{L}_{\text{DICE}}(\tilde{\mathbf{y}}_{\text{mask}},\mathbf{y}_{\text{mask}}),$
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are the weight of the above losses, following LISA [24] set to 1.0, 2.0, and 0.5, respectively.
In the pretraining phase, we mainly adopt the mixed dataset configurations in LISA [24] to mix four types of datasets and sample in training splits by the ratio 9 of Semantic Segmentation (SemSeg), 6 of Referring Expression Segmentation (RES), 3 of Visual Question Answering (VQA), and 1 of Reasoning Segmentation (ReasonSeg). This ratio of 9:6:3:1 is almost kept from LISA for a fair comparison, except for the increment of RES to weight gRefCOCO [30] more. For each task, the sample is uniformly sampled across the dataset. The semantic segmentation datasets consist of ADE20K [73], COCO-Stuff [3], Mapillary Vistas [38], PACO-LVIS [46], and PASCAL-Part [5]. The RES datasets include RefCOCO, RefCOCO+ [22], RefCOCOg [37], RefCLEF [22], and gRefCOCO [30]. VQA and reasoning segmentation adopt LLaVA-Instruct-150K [31] and ReasonSeg [24] as the datasets, respectively.
We train all three model variants (Vicuna-7B, Vicuna-13B, and Llama2-13B) on 8 Tesla A100 GPUs (80GB) for 50,000 steps with a batch size of 2 per device. The learning rate is set to $3\!\times{}\!10^{-4}$ with a gradient accumulation of 10. The weight decay is set to 0, and the gradients are clipped to 1 by the maximum norm. The learning rate is warmed up in 100 steps and linearly decayed. To reduce the GPU memory footprints, we adopt the AdamW [34] optimizer with stage 2 ZeRO [45]. As for LoRA, we pick 8 as the LoRA rank for the 7B variant and 64 for the 13B variants.
Finetuning on GRES task. We finetune the pretrained models on gRefCOCO [30] dataset to improve the GRES performances further. For all variants of GSVA, we load the weights of the respective model in pretraining and train another 10 epochs on gRefCOCO. This is slightly different from pretraining because we go through the whole training set once per epoch without sampling data from it. In the finetuning experiments, we use a lower learning rate $1\!\times{}\!10^{-4}$, keeping other configurations unchanged.
Finetuning on classic RES task. For classic RES, including RefCOCO, RefCOCO+ [22], and RefCOCOg [37], we choose the mixed classic RES dataset during training aside from gRefCOCO. In addition to reducing the learning rate to $1\!\times{}\!10^{-4}$, the other hyper-parameters remain unchanged.
Appendix B GSVA on Ref-ZOM
Ref-ZOM dataset. Proposed by DMMI [19], Ref-ZOM is a similar dataset to gRefCOCO [30], posing the challenges of one-to-one, one-to-many, and one-to-zero referring expression to targets in the image relationships in RES. The one-to-one is the case of classic RES where the referring expression is matched with only one target in the image, while the one-to-many and one-to-zero are the multi-target and empty-target cases, respectively. Ref-ZOM contains 55,075 images and 74,942 annotated objects, among which there are 56,972 one-to-one cases, 21,290 one-to-many cases, and 11,937 one-to-zero cases. DMMI provides a default split of training set and test set of Ref-ZOM. There are 43,749 images in the training set and 11,329 images in the test set.
Setups. We regard Ref-ZOM [19] as a kind of GRES, and use a similar protocol to gRefCOCO [30] to evaluate LISA and GSVA on Ref-ZOM dataset. For Ref-ZOM, gIoU and cIoU are substituted to the equivalent metrics, mIoU and oIoU. Different from gRefCOCO, mIoU, and oIoU only count for non-empty targets. For one-to-zero, i.e., empty targets, the accuracy is one if the predicted mask is strictly all-zero. LISA [24] and our proposed GSVA with Vicuna-7B [8] are evaluated, including the pretrained versions and the finetuned versions with 1 epoch on the training split.
Results. As shown in Table 6, without finetuning on Ref-ZOM, GSVA surpasses LISA by clear margins in oIoU and mIoU of over 6%. Also, GSVA approaches close to the SOTA, DMMI [19], which is the SOTA proposed along with the Ref-ZOM dataset. After finetuning for 1 epoch, LISA quickly catches up GSVA with less 2% IoU metrics, while GSVA achieves 94.59% accuracy in classifying empty targets, keeping a competitive performance to DMMI.
Method Ref-ZOM [19] Test Set
oIoU mIoU Acc
MCN [35] 55.03 54.70 75.81
CMPC [20] 56.19 55.72 77.01
VLT [11] 60.21 60.43 79.26
LAVT [62] 64.45 64.78 83.11
DMMI [19] 68.77 68.21 87.02
LISA-Vicuna-7B [24] 60.14 61.46 72.58
GSVA-Vicuna-7B 67.12 67.98 82.66
LISA-Vicuna-7B [24] (ft) 66.41 65.39 93.39
GSVA-Vicuna-7B (ft) 68.29 68.13 94.59
Table 6: GRES results on the test split of Ref-ZOM [19] dataset. oIoU and mIoU are only computed on the samples containing targets, while the correct prediction of Acc is the mask of all zeros for empty targets. Baselines are excerpted from  Hu et al. [19].
Appendix C GSVA on Semantic Segmentation
To verify the vanilla semantic segmentation ability, we include an extra evaluation of the pretrained GSVA-Vicuna-7B and LISA-Vicuna-7B on the ADE20K [73] validation dataset. Since the pretraining covers the training set of ADE20K, we only evaluate them to see if they are capable of segmenting semantic regions rather than some specific objects. We prompt the models with the instruction as User: What is {classname} in this image? Assistant: Sure, [SEG]., where each {classname} is filled with the class name that exists in this image. We report the mIoU of the predictions and the ground truths, in Table 7. LISA achieves 60.11 mIoU whereas our GSVA slightly improves to 60.56.
Model mIoU(ss)
LISA-Vicuna-7B [24] 60.11
GSVA-Vicuna-7B 60.56
Table 7: Semantic Segmentation Results of GSVA and LISA [24] on ADE20K validation dataset. The mIoU is different from the one in the closed-set segmentation, which is computed with the existing classes, and resembles the recall to some extent.
Appendix D More Ablation Study
Figure 5: Ablation on how many [SEG] tokens are used during training. Metrics including gIoU, cIoU, and N-acc on gRefCOCO [30] dataset are reported.
The number of targets in training. We study the effect on how many [SEG] tokens are adopted in training GSVA. In GSVA, the number of [SEG]s controls the number of targets processed by GSVA, which impacts the final capacity of multiple target segmentation. We sweep this hyper-parameter from 1 to 10 and report the results of GSVA-Vicuna-7B on gRefCOCO [30] validation set in Figure 5. On the one hand, from the figure, all the metrics include gIoU, cIoU, and N-acc. increases as the number of [SEG] grows from 1 to 5, and the model performances saturate after 5 [SEG] tokens are involved. On the other hand, the memory footprints keep rising as more [SEG] tokens are used and exceed 24GB (RTX3090 / RTX4090) after 6 [SEG] tokens. Considering the performance saturation and the memory consumption, we pick 5 as the default.
Weight sharing for different [SEG] tokens. Another design choice considering multiple [SEG] tokens is to use a set of individual [SEG] embeddings rather than sharing weights in GSVA. We adopt 8 tokens from [SEG000] to [SEG007] to study this design. In case of more than 8 targets, we dispatch the targets based on the remainder after divided by 8. However, this design poses convergence difficulties, as shown in Table 8. We attribute the intensive performance degradation to the slow convergence of the [SEG] token embedding, which is that too many [SEG] token embeddings could interfere with each other. Therefore it is reasonable to share weights between [SEG] tokens.
Model gRefCOCO Val.
gIoU cIoU N-acc.
GSVA (weight sharing) 63.32 61.70 56.45
w/ 8 individual embeddings 16.13 23.11 0.00
Table 8: Ablation study on weight sharing of [SEG] tokens.
Different SAM Backbones. We alter the default pretrained SAM-ViT-H [23] vision encoder to SAM-ViT-L and SAM-ViT-B, to verify the effectiveness of GSVA that is stable to the change of segmentation vision encoder $F_{\text{V2}}$. As shown in Table 9, when we substitute the ViT-H [13] backbones with SAM pretrained ViT-B/L, the segmentation performances decreases by 4$\sim$5while the accuracy of empty targets change mildly and even increases to 58.57% with ViT-B. The slight drops of gIoU and cIoU scores could be attributed to the usages of more miniature vision encoders, and the stable, even higher N-acc implies that generating [REJ] token by MLLM consistently works well with SFMs in different capacities. These results also demonstrate the effectiveness of the design that rejects the empty targets through a special token rather than producing an all-zero mask.
Vision Encoder $F_{\text{V2}}$ in SFM gRefCOCO Val.
gIoU cIoU N-acc.
SAM-ViT-H (GSVA) 63.32 61.70 56.45
SAM-ViT-L 61.67 60.46 56.04
SAM-ViT-B 59.20 56.36 58.57
Table 9: Ablation study on Different SAM [23] ViT backbones.
Appendix E More Visualizations
We provide more visualization results in Figure 6 to further show the effectiveness of our GSVA in GRES task.
(a) Examples of GSVA on gRefCOCO [30] test set A and B. (I)
(b) Examples of GSVA on gRefCOCO [30] test set A and B. (II)
(c) Examples of GSVA on gRefCOCO [30] test set A and B. (III)
(d) Examples of GSVA on gRefCOCO [30] test set A and B. (IV)
(e) Examples of GSVA on gRefCOCO [30] test set A and B. (V)
(f) Examples of GSVA on gRefCOCO [30] test set A and B. (VI)
Figure 6: Visualization of GSVA in GRES task, the inputs and outputs are presented in the form of dialogues between human user and the chatbot. The examples are selected from gRefCOCO [30] test set A and B. Zoom in for best view.